OpenStudy (anonymous):

write a slope intercept equation of the line whose graph is described Parallel to the graph of 3x + 2y = 5 y-intercept (0,-1)

OpenStudy (anonymous):

First the slope: 3x + 2y = 5 2y = -3x + 5 y = -3/2x + 5/2 slope = -3/2 Intercept = -1 equation y = -3/2x -1